OpenStudy (anonymous):

Solve the following equation for n. P/3 +A= 4m-3n n=?

OpenStudy (blurbendy):

Start by getting 'n' by itself (hint: move everything to the left side).

OpenStudy (anonymous):

So then it's p/3+A-4m= -3n?

OpenStudy (blurbendy):

you'e almost there. we still have to deal with that -3 in front of the n. do you have any ideas how we could move that - 3 to the left side?

OpenStudy (anonymous):

p/3+a-4m/-3=n?

OpenStudy (blurbendy):

yes, EVERYTHING on the left side is divided by -3 good job

OpenStudy (anonymous):

So n= 1/9(-3A+12m-P)?

OpenStudy (blurbendy):

perfect

OpenStudy (anonymous):

So that's the answer?

OpenStudy (blurbendy):

looks good to me